OpenStudy (anonymous):

find the point slope form of an equation of the line thru the two ordered pairs (7,15)(12,92)

OpenStudy (anonymous):

y-15=(slope)(x-7) where your slope is (92-15)/(12-7) i.e 77/5 y-15=(77/5)(x-7) 7y-105=77x-329 77x-7y-224=0

OpenStudy (anonymous):

means 11x-y-32=0